OpenStudy (anonymous):

int x/sqrt(x+1)dx please assist :)

OpenStudy (anonymous):

sorry, missed the sqrt part. just a sec.

OpenStudy (anonymous):

i think this is a trig substitution and i'm so bad at those lol

ganeshie8 (ganeshie8):

try x+1 = u^2

OpenStudy (anonymous):

\[\int\limits \frac{ x }{ \sqrt{x+1} }dx=\int\limits x \left( x+1 \right)^{-\frac{ 1 }{ 2 }}\,dx\] you could do integration by parts... have you learned that yet?

OpenStudy (anonymous):

yeah, would it be u=(x+1)^(1/2), dv=xdx

OpenStudy (anonymous):

no, let u = x

OpenStudy (anonymous):

then du = dx and you get rid of the "bad" part. you can integrate 1/(x+1)

OpenStudy (anonymous):

you missed the sqrt again lol

OpenStudy (anonymous):

sorry, that doesn't prevent it from being integrable, though

OpenStudy (anonymous):

ok so do you end up with

OpenStudy (anonymous):

u/sqrt(x+1) du

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{ \sqrt{x+1} }=\int\limits u^{-\frac{ 1 }{ 2 }}du=2u^{\frac{ 1 }{ 2 }}+c\]

OpenStudy (anonymous):

\[=2\sqrt{x+1}\]

OpenStudy (anonymous):

+c

OpenStudy (anonymous):

but you just lost the xdx and replaced it with just dx how did that happen

OpenStudy (anonymous):